OpenStudy (anonymous):

You are remodeling your kitchen. You've contacted two tiling companies who gladly told you how long it took their workers to tile a floor of a similar size. Jim completed half the floor in 12 hours. Pete completed the other half of the floor in 10 hours. If Pete can lay 25 more tiles per hour than Jim, what equation would you use to find the rate that Jim can lay tiles? Let x represent Jim's rate.

OpenStudy (anonymous):

12x=10(x-25) 12x=10(x+25) 10x=12(x-25) 10x=12(x+25)

OpenStudy (anonymous):

I guess B

OpenStudy (anonymous):

help?